Toss a fair coin twice. You win Ghc 1 if at least one of the two tosses comes out heads.

(a) Assume that you play this game 300 times. What is, approximately, the probability that you win at least Ghc 250?

(b) Approximately how many times do you need to play so that you win at least Ghc 250 with probability at least 0.99?



1
Expert's answer
2022-04-27T02:12:25-0400

P(winning Ghc 1) = P(getting at least 1 heads in two tosses)"=\frac34"

Now, n=300, p=3/4, q=1/4

"\mu=np=225=\sqrt{56.25}=7.5"

"\sigma=\sqrt{npq}"

X∼N(μ,σ)

a. "P(X\u2265250)=1\u2212P(X<250)=1\u2212P(z< \n\\frac{\n250\u2212225\n\u200b}{7.5}\n )\n\n=1-P(z<3.3)=1-0.99957=0.00043"

b. "P(X\ge250)\ge0.99 \Rightarrow 1-P(X<250)\ge0.99 \\ \Rightarrow 0.01\ge P(X<250) \\ \Rightarrow P(z<\frac{250-n(3/4)}{\sqrt{n(3/4)(1/4)}})\le 0.01
\Rightarrow \frac{250-n(3/4)}{\sqrt{n(3/4)(1/4)}}>0.50399 \\ \Rightarrow \frac{250-n(3/4)}{0.50399}>\sqrt{(3n/16)} \\ \Rightarrow ( \frac{250-n(3/4)}{0.50399})^2>3n/16"

On solving, we get,


n>338.68


Or "n\ge339"